# The Rayleigh-Taylor instability with foams

### v1

@@ Abstract
We analyse the behaviour of the Rayleigh-Taylor instability (RTI) in the presence of a foam. Such a problem may be relevant, for example, to some inertial confinement fusion (ICF) scenarios such as foams within the capsule or lining the inner hohlraum wall. The foam displays 3 different phases: by order of increasing stress, it is first elastic, then plastic, and then fractures. Only the elastic and plastic phases can be subject to a linear analysis of the instability. The growth rate is analytically computed in these 2 phases, in terms of the micro-structure of the foam. In the first, elastic, phase, the RTI can be stabilized for some wavelengths. In this elastic phase, a homogenous foam model overestimates the growth because it ignores the elastic nature of the foam. Although this result is derived for a simplified foam model, it is likely valid for most of them. Besides the ICF context considered here, our results could be relevant for many fields of science.
@@ I Introduction
With ignition reached several times at Livermore, inertial confinement fusion enters a new era where the goal is clearly to increase the yield and the repetition rate Zylstra et al. (2022); Abu-Shawareb et al. (2022, 2024). In this respect, the use of foams in the target has been contemplated for some time by some authors as a means to increase laser-target coupling and to more easily and cheaply mass-produce targets compared to what is possible with solid ice layered ones Haines (1997); Depierreux et al. (2009); Goncharov et al. (2020); Paddock et al. (2023).
@@
In parallel, it has been recognized for long that a paramount process during the target implosion is the Rayleigh-Taylor Instability (RTI - See Hurricane et al. (2023) and references therein). In this respect, the question surges immediately: how does the RTI behave when a foam is involved? At one end of the theoretical spectrum, one can answer the question ignoring the microstructure of the foam, considering it a homogeneous medium of a designated average density. At the other end of the same spectrum, the behaviour of the RTI when an intact foam is involved, is an open question.
The present work does not aim at filling the theoretical gap between intact and homogenised foam, but at exploring the “intact” end of the gap. Namely, how does the RTI behaves when an intact foam is involved? To which extent does it differ from that of a homogenous fluid?
Notably, a foam can be “dry” or “wetted”. The latter would be comparable to a wetted sponge. Even though wetted foams are more relevant to ICF than dry ones Paddock et al. (2023), we shall here consider dry foams. To our knowledge, there is currently no theory of the mechanical properties of wetted foams, while there is for dry foams. This is why the “wetted counterpart” of the present work is left for future works.
@@
Our findings are summarized in the conclusion, where we also explain why the reduction of the RTI growth rate in the elastic phase is likely valid for most foams, even though the present work focuses on a simplified model.
@@ II Model of foam
Foams come with a great variety of flavour, like two-dimensional honeycomb or three-dimensional foams, with open or closed cells, etc. We shall here focus on a two-dimensional honeycomb, the cell unit of which is represented in Fig. 1.
@@
Figure 2: Typical stress-strain curve of a foam. See Eq. (12) for $\varepsilon_{D}$. Adapted from Gibson et al. (1982).
@@ II.1 Elastic phase
For a regular pattern with $\theta=30^{\circ}$ and $h=l$, the stress tensor is isotropic. $E^{*}_{1}$ and $E^{*}_{2}$ then reduce to,
@@ II.2 Plastic phase
Putting together Eqs. (6,8), we can derive the elastic collapse strain $\varepsilon^{*}_{el}$ corresponding to such a stress,
@@ II.3 Fracture phase
We shall model the green plateau by a horizontal line, implying, from Eqs. (8, 10)
$(\sigma^{*}_{el})_{2}$ $=$ $\sigma^{*}_{pl}$
@@ III RTI formalism
Consider the setup pictured in Figure 3. The foam average density is $\rho$. It is placed below a fluid of density $\rho_{up}>\rho$. At equilibrium, the pressure at the interface of the two media is $P_{0}$. The interface is now bent over a distance $\sim 1/k\ll L$, by an amplitude $\delta X_{2}$. We assume the foam uniformly supports the higher density fluid above it, so that the pore size $\sim 2l\cos\theta=l\sqrt{3}$ (see Fig. 1 for $\theta=30^{\circ}$) is smaller than the wavelength of the perturbation, namely
@@
$=$ $(\rho_{up}-\rho)~g\delta X_{2}S$
$=$ $(\rho_{up}-\rho)~g\delta X_{2}\frac{D}{k}.$
@@
$(\rho_{up}+\rho)\frac{D}{k^{2}}\ddot{\delta X_{2}}=(\rho_{up}-\rho)~g\delta X_{2}\frac{D}{k},$ (17)
@@ IV RTI with a foam
Are the three phases of the foam behavior amenable to a linear RTI analysis? No.
The linear theory of the RTI requires $\delta X_{2}\ll 1/k$, that is, $\varepsilon\equiv k\times\delta X_{2}\ll 1$. Figure 2, with $\varepsilon^{*}_{el}$ defined by Eq. (II.2), shows that the full elastic phase of the foam definitely fits into the linear regime. Part of the green plateau phase equally fulfills the linear requirement since it starts from $\varepsilon\sim 10^{-1}$, as evidenced by Eq. (II.2). Yet, the last phase, the fracture phase of the foam, evidenced by a nearly vertical line on Figure 2, starts, according to Eq. (13), from $\varepsilon_{D}=1-1.61t/l\sim 1^{-}$. The foam will therefore leave the linear regime somewhere along the plateau, rendering the linear analysis of the last phase invalid.
@@ IV.1 Elastic phase
$=$ $(\rho_{up}-\rho)~g\delta X_{2}S-Ek\delta X_{2}S$
$=$ $(\rho_{up}-\rho)~g\delta X_{2}\frac{D}{k}-Ek\delta X_{2}\frac{D}{k}$
@@ IV.2 Plastic phase
$=$ $(\rho_{up}-\rho)~g\delta X_{2}S-\sigma_{el}^{*}S$
$=$ $(\rho_{up}-\rho)~g\delta X_{2}\frac{D}{k}-\sigma_{el}^{*}\frac{D}{k},$
@@
$(\rho_{up}+\rho)\frac{D}{k^{2}}\ddot{\delta X_{2}}=(\rho_{up}-\rho)~g\delta X_{2}\frac{D}{k}-\sigma_{el}^{*}\frac{D}{k},$ (29)
@@ V Conclusion
We can now assess the difference between the RTI with the foam and with the equivalent medium of average density $\rho$. Such a difference is only notable in the first phase, namely the elastic phase, since the growth rate of the RTI in the next phase, the plastic one, is the same in both cases (see Section IV.2).
From Eqs. (19,24), we can express the ratio of the growth rate $\gamma^{\prime}$ of the foam-RTI, to the growth rate $\gamma$ of the averaged medium-RTI,
@@
After replacing $\rho$ and $E$ in the expression (25) of $k_{m}$ by Eqs. (2,6), we find an expression of $k_{m}$ in terms of the acceleration $g$, the density $\rho_{up}$, and the properties of the material the foam is made of,
@@
The present results have been derived for a simplified model of foam. Yet, all foams seem to exhibit an elastic phase at the beginning of their strain-stress curve, be it in 2D Gibson et al. (1982) or even in 3D Gibson and Ashby (1982). Since our key result, namely the reduction of the RTI growth rate in the elastic phase, relies on the existence of such a phase, our conclusion is likely valid for most foams. Only the Young modulus $E$ involved in Figure 2 and Section IV.1 needs to be adapted to the specific foam under scrutiny.

### v2

# The linear Rayleigh-Taylor instability with foams

@@ Abstract
We analyse the behaviour of the linear phase of the Rayleigh-Taylor instability (RTI) in the presence of a foam. Such a problem may be relevant, for example, to some inertial confinement fusion (ICF) scenarios such as foams within the capsule or lining the inner hohlraum wall. The foam displays 3 different phases: by order of increasing stress, it is first elastic, then plastic, and then fractures. Only the elastic and plastic phases can be subject to a linear analysis of the instability. The growth rate is analytically computed in these 2 phases, in terms of the micro-structure of the foam. In the first, elastic, phase, the RTI can be stabilized for some wavelengths. In this elastic phase, a homogenous foam model overestimates the growth because it ignores the elastic nature of the foam. Although this result is derived for a simplified foam model, it is likely valid for most of them. Besides the ICF context considered here, our results could be relevant to many fields of science.
@@ I Introduction
With ignition reached several times at Livermore, inertial confinement fusion enters a new era where the goal is clearly to increase the yield and the repetition rate Zylstra et al. (2022); Abu-Shawareb et al. (2022, 2024). In this respect, the use of foams in the target has been contemplated for some time by some authors as a means to increase laser-target coupling and to more easily and cheaply mass-produce targets compared to what is possible with solid ice layered ones Haines (1997); Depierreux et al. (2009b); Goncharov et al. (2020); Paddock et al. (2023). Hence, in connection to ICF, foams have been studied theoretically Gus’kov and Rozanov (1997); Gus’kov et al. (2011, 2015); Belyaev et al. (2018), numerically Kapin et al. (2006); Velechovsky et al. (2016); Belyaev et al. (2020); Milovich et al. (2021); Hudec et al. (2023) and experimentally Limpouch et al. (2006, 2008); Depierreux et al. (2009a); Nicolaï et al. (2012); Cipriani et al. (2018); Tikhonchuk et al. (2019); Cipriani et al. (2021); Igumenshchev et al. (2023).
@@
In parallel, it has been recognized for long that a paramount process during the target implosion is the Rayleigh-Taylor Instability (RTI - See Hurricane et al. (2023) and references therein). In this respect, the question surges immediately: how does the RTI behave when a foam is involved? In ICF experiments, the foam is anticipated to ionize and homogenize rapidly, and so at one end of the theoretical spectrum, one can answer the question ignoring the microstructure of the foam, considering it a homogeneous medium of a designated average density. At the other end of the same spectrum, the behaviour of the RTI when an intact foam is involved, is an open question that we here address.
Experimental evidence of foam pore structure spatial imprint on transmitted shock fronts has recently been observed. In planar shock uniformity experiments conducted at OMEGA upon wetted foam samples, imaging of transmitted shock fronts with the Omega High Resolution Velocimeter revealed a spatial imprint of various pore structures upon the shock front after breakout from the back of the foam material DeVault et al. (2026). This observation suggests that the discrete structure of foams does influence shock front progression, making clear the necessity for the development of models which describe the impact of foam pore structures on shock front progression and instability growth.
The present work does not aim at filling the theoretical gap between intact and homogenised foam, but at exploring the “intact” end of the gap. Namely, how does the RTI behaves when an intact foam is involved? To which extent does it differ from that of a homogenous fluid? Though an intact foam framework is not expected to remain valid throughout a laser-driven ICF implosion, this work defines the limiting behavior of intact foams and explores to what extent it differs from that of a homogeneous fluid, providing the two end points of the foam-RTI spectrum.
Clearly, the foam will certainly not be intact in ICF conditions. Yet, the “imprint experiment” DeVault et al. (2026) shows that the foam may not be 100% homogenized either. We therefore think that in order to investigate the behavior of the foam in real conditions, elucidating the other extreme state, namely the intact state, may prove useful.
Notably, a foam can be “dry” or “wetted”. The latter would be comparable to a wetted sponge. Even though wetted foams are more relevant to ICF than dry ones Paddock et al. (2023), we shall here consider dry foams, which, for instance, have been used to line the inner wall of hohlraums at the NIF as a possible approach to tamping hohlraum wall expansion and improving drive symmetry control Moore et al. (2020b, a); Bhandarkar et al. (2018).
To our knowledge, there is currently no theory of the mechanical properties of wetted foams, while there is for dry foams. This is why the “wetted counterpart” of the present work is left for future works.
@@
Our main finding is as follows: as they are deformed, foams first exhibit an elastic phase with a spring-like behaviour. Then, they go through a plastic phase (see Figure 2). In the elastic phase, the growth of the RTI is reduced, and even suppressed beyond a critical $k$ given by Eq. (27). Analytical results are presented, for a specific model of foam, in terms of the foam inner geometry.
The taming of the RTI in elastic materials is not new Dimonte et al. (1998); Piriz et al. (2005, 2009b, 2009a). What we think is new is that such a taming applies to foams. And likely to all of them, for all of them share the same elastic behavior for small strain Gibson et al. (1982); Gibson and Ashby (1982).
@@ II Model of foam
Foams come with a great variety of flavour, like two-dimensional honeycomb or three-dimensional foams, with open or closed cells, etc. We shall here focus on a two-dimensional honeycomb, the cell unit of which is represented in Figure 1.
@@
Figure 2: Typical stress-strain curve of a foam. See Eq. (II.2) for $\varepsilon^{*}_{el}$ and Eq. (13) for $\varepsilon_{D}$. The value of $E$, slope of the red line, is given by Eq. (7). Adapted from Gibson et al. (1982).
@@ II.1 Elastic phase
For a regular pattern with
$\theta$ $=$ $30^{\circ},$ (6)
$h$ $=$ $l,$
the stress tensor is isotropic. $E^{*}_{1}$ and $E^{*}_{2}$ then reduce to,
@@ II.2 Plastic phase
Putting together Eqs. (7,9), we can derive the elastic collapse strain $\varepsilon^{*}_{el}$ corresponding to such a stress,
@@ II.3 Fracture phase
We shall model the green plateau by a horizontal line, implying, from Eqs. (9, 11)
$\sigma^{*}_{el}$ $=$ $\sigma^{*}_{pl}$
@@ III RTI formalism
Consider the setup pictured in Figure 3. The foam average density is $\rho$. It is placed below a fluid of density $\rho_{up}>\rho$. At equilibrium, the pressure at the interface of the two media is $P_{0}$. The interface is now bent over a distance $\sim 1/k\ll L$, by an amplitude $\delta X_{2}$. We assume the foam uniformly supports the higher density fluid above it, so that the pore size $\sim 2l\cos\theta=l\sqrt{3}$ (see Figure 1 for $\theta=30^{\circ}$) is smaller than the wavelength of the perturbation, namely
@@
$=$ $(\rho_{up}-\rho)~g~\delta X_{2}S$
$=$ $(\rho_{up}-\rho)~g~\delta X_{2}\frac{D}{k}.$
@@
$(\rho_{up}+\rho)\frac{D}{k^{2}}\ddot{\delta X_{2}}=(\rho_{up}-\rho)~g~\delta X_{2}\frac{D}{k},$ (18)
@@ IV RTI with a foam
IV.1 Limits of the linear analysis
Are the three phases of the foam behavior amenable to a linear RTI analysis?
Unstable RTI modes grow exponentially until their amplitude reaches an $\alpha$ fraction of their wavelength $\lambda=2\pi/k$ Haan (1989); Zhou (2017a, b), with $\alpha\in[1/10,1/5]$. Therefore, the linear regime holds until $\delta X_{2}=\alpha\lambda$, that is,
$k\times\delta X_{2}=\varepsilon=2\pi\alpha\in[0.6,1.2].$ (23)
Figure 2, with $\varepsilon^{*}_{el}\sim 0.1$ defined by Eq. (II.2), shows that the full elastic phase of the foam fits into the linear regime, even with $\alpha=1/10$. Part of the green plateau phase should equally fulfills the linear requirement (23).
Regarding the last phase, the fracture phase of the foam, evidenced by a nearly vertical line on Figure 2, Eq. (14) shows it starts from $\varepsilon_{D}=1-1.61t/l\sim 1^{-}$.
Therefore, whether the full plateau and the fracture phases fit the linear regime, depends on $\alpha$. Future numerical simulation could help narrowing down its value for the present scenario.
@@ IV.2 Elastic phase
$=$ $(\rho_{up}-\rho)~g~\delta X_{2}S-Ek~\delta X_{2}S$
$=$ $(\rho_{up}-\rho)~g~\delta X_{2}\frac{D}{k}-Ek~\delta X_{2}\frac{D}{k}$
@@ IV.3 Plastic phase
$=$ $(\rho_{up}-\rho)~g~\delta X_{2}S-\sigma_{el}^{*}S$
$=$ $(\rho_{up}-\rho)~g~\delta X_{2}\frac{D}{k}-\sigma_{el}^{*}\frac{D}{k},$
@@
$(\rho_{up}+\rho)\frac{D}{k^{2}}\ddot{\delta X_{2}}=(\rho_{up}-\rho)~g~\delta X_{2}\frac{D}{k}-\sigma_{el}^{*}\frac{D}{k},$ (31)
@@
Figure 4: Plot of Eq. (33) for 3 values of $\theta$ for $n=0.69$.
IV.4 Foam design for RTI mitigation
We here comment on the possibility of tailoring the foam design for applications where control of instability is important.
Since the taming of the RTI is found in the foam elastic phase, it could be beneficial for stability purposes to extend it as far as possible in terms of the strain $\varepsilon$. For the model of foam chosen here, namely $\theta=30^{\circ}$, $h=l$ and $n=0.69$, $\varepsilon^{*}_{el}$ is set to $\sim 0.1$, as fixed by Eq. (II.2). However, coming back to the general expressions (5,8) of the slope of the red line on Figure 2, and of the critical stress $\sigma^{*}_{el}$ marking the onset of the plateau, we can obtain a more general expression of $\varepsilon^{*}_{el}$,
$\varepsilon^{*}_{el}=\frac{\pi^{2}n^{2}}{24}\frac{(l/h)^{3}\cos^{2}(\theta)}{(l/h)\sin(\theta)+1}.$ (33)
Figure 4 plots this equation for $n=0.69$ and 3 values of $\theta$. The higher $l/h$, the longer the elastic phase lasts. Simply put, cells foams should be wider than high.
Varying $\varepsilon^{*}_{el}$ by controlling $l/h$ could help test the theory by measuring growth rates, seeing if, as predicted, an abrupt change occurs at the elastic-plastic transition.
The maximum growth rate (29) in the elastic phase can can also be tailored. It reads
$\gamma^{\prime 2}=g^{2}A\frac{\rho_{up}-\rho}{4E}=g^{2}\frac{(\rho_{up}-\rho)^{2}}{\rho_{up}+\rho}\frac{1}{4E}.$ (34)
Here, the inner foam design comes into play through Eqs. (1,5), defining respectively the average density $\rho$ and the Young modulus $E$ (now in the $X_{2}$ direction). Contrary to Eq. (33), this expression involves too many parameters to be simply analyzed. However, for large values of $l/h$ (and $\theta\neq 0$), $\varepsilon^{*}_{el}$ given by (33) scales like $(l/h)^{2}$. This allows to replace $l/h$ by its expression in terms of $\varepsilon^{*}_{el}$ in the equation (34) above. The leading term of the result for large $l/h$, reads,
$\gamma^{\prime 2}\sim g^{2}\frac{12\sqrt{6}}{\pi^{3}n^{3}}\frac{\rho_{up}}{E_{s}}\left(\frac{h}{t}\right)^{3}\sqrt{\sin\theta}~\varepsilon^{*3/2}_{el}.$ (35)
It seems then that a large $\varepsilon^{*}_{el}$, that is, a long elastic phase, also favors a higher growth rate. A detailed analysis including all the factors involved is required to assess the trade off between both quantities. Note also that for foam parameters different than those defined by Eq. (6), the Young moduli in the $X_{1}$ and $X_{2}$ will be different. The anisotropy so introduced should modify the RTI analysis of Sections III & IV.
@@ V Conclusion
We can now assess the difference between the RTI with the foam and with the equivalent medium of average density $\rho$. Such a difference is only notable in the first phase, namely the elastic phase, since the growth rate of the RTI in the next phase, the plastic one, is the same in both cases (see Section IV.3).
From Eqs. (20,26), we can express the ratio of the growth rate $\gamma^{\prime}$ of the foam-RTI, to the growth rate $\gamma$ of the averaged medium-RTI,
@@
After replacing $\rho$ and $E$ in the expression (27) of $k_{m}$ by Eqs. (2,7), we find an expression of $k_{m}$ in terms of the acceleration $g$, the density $\rho_{up}$, and the properties of the material the foam is made of,
@@
The present results have been derived for a simplified model of foam. Yet, all foams seem to exhibit an elastic phase at the beginning of their strain-stress curve, be it in 2D Gibson et al. (1982) or even in 3D Gibson and Ashby (1982). Since our key result, namely the reduction of the RTI growth rate in the elastic phase, relies on the existence of such a phase, our conclusion is likely valid for most foams. Only the Young modulus $E$ involved in Figure 2 and Section IV.2 needs to be adapted to the specific foam under scrutiny.